Documenta Mathematica

Band 16, 2011

Richard Hill and David Loeffler Emerton’s Jacquet Functors for

Non-Borel Parabolic Subgroups  1–31

Chiarellotto, B. and Tsuzuki, N. Log-Growth Filtration and Frobenius Slope Filtration of F -Isocrystals

at the Generic and Special Points  33–69

Jürgen Hausen, Elaine Herppich, Hendrik Süß Multigraded Factorial Rings

and Fano Varieties with Torus Action  71–109

Rune Johansen

On the Structure of Covers of Sofic Shifts  111–131

Kengo Matsumoto Ergodic Properties and KMS Conditions

on C^{∗}-Symbolic Dynamical Systems  133–175

M.Koenenberg

An Infinite Level Atom coupled to a Heat Bath  177–208

Markus Perling

Divisorial Cohomology Vanishing on Toric Varieties  209–251

Richard Pink, Torsten Wedhorn, and Paul Ziegler

Algebraic Zip Data  253–300

Thomas M. Fiore, Wolfgang Lück, and Roman Sauer Euler Characteristics of Categories

and Homotopy Colimits  301–354

L. Battisti Surfaces de Stein Associées

aux Surfaces de Kato Intermédiaires  355–371

G. A. Noskov Asymptotic Behavior

of Word Metrics on Coxeter Groups  373–398

Ugo Bruzzo and Dimitri Markushevish

Moduli of Framed Sheaves on Projective Surfaces  399–410

F. Déglise

Modules Homotopiques  411–455

Daniel R. Grayson

The Additivity Theorem in Algebraic K-Theory  457–464

Marcus Zibrowius

Witt Groups of Complex Cellular Varieties  465–511

Joseph Ross The Hilbert-Chow Morphism

and the Incidence Divisor  513–543

Laurent Clozel Presentation of an Iwasawa Algebra:

The Case of Γ_{1}SL(2, ℤ_{p})  545–559

Vitali Vougalter, Vitaly Volpert On the Existence of Stationary Solutions for

Some Non-Fredholm Integro-Differential Equations  561–580

Claude Cibils, María Julia Redondo, and Andrea Solotar Fundamental Group of Schurian Categories

and the Hurewicz Isomorphism  581–595

Douglas S. Bridges Characterising Weak-Operator

Continuous Linear Functionals on B(H) constructively 597–617

Jeffrey L. Boersema, Efren Ruiz, P. J. Stacey The Classification of Real

Purely Infinite Simple C*-Algebras  619–655

Gonçalo Tabuada A Simple Criterion for Extending

Natural Transformations to Higher K-Theory  657–668